How many ways are there to choose a committee of 7 people from a group of 10 people?
Found 2 solutions by stanbon, thejackal:
Answer by stanbon(75887)

How many ways are there to choose a committee of 7 people from a group of 10 people?
-------
Ans: 10C7 = 10C3 = (10*9*8)/(1*2*3) = 720/6 = 120
---------------
Cheers,
Stan H.
-------------------

Answer by thejackal(72)

lets write it differently shall we.
there are 10 committee members, so lets number them as committee member numbers 1,2,3,4,5,6,7,8,9,10
we want seven. First of all there are a few very important things to note about the question.
1. Order is not important. I need to choose 7 members, not 7 consecutive members but any 7 from 10. From that we can deduce that this is a combinatorics problem of type n choose k where n = size of the larger set and k = the size of each combination
n choose k is found by calculating the binomial coefficient found using the formula
n!/k!(n-k)! thus our answer is 10!/7!(10-7)! = 120 combinations of 7 committee members
ref: http://en.wikipedia.org/wiki/Combination
useful link: combinatorics and permutation calculator -http://www.mathsisfun.com/combinatorics/combinations-permutations-calculator.html